A piano has been pushed to the top of the ramp at the back of a moving van. The workers think it is safe, but as they walk away, it begins to roll down the ramp. Neglect the friction between the piano and the ramp.

If the back of the truck is 1.5 m above the ground and the ramp is inclined at 22 ∘ , how much time do the workers have to get to the piano before it reaches the bottom of the ramp?
Answer #1

answer) the force acting in vertical direction

Fy=m*g* sin........1)

by newtons law we have

Fy=ma

so the eqn 1 becomes

ma=mgsin

a=gsin

a=9.8*sin22=3.67114 m/s2

using kinematics we have the formula

s=1.5/sin22=4.0042 m

s=1/2at2

4.0042=1/2*3.67114*t2

t=2.1814=1.5 s

so the answer is 1.5 s or 1.48 s